OpenStudy (anonymous):

(e^x)e^(x+1)=1 in logarithmic equation, solve the equation

OpenStudy (anonymous):

(e^x)e^(x+1) = 1 e^(2x + 1) = 1 Take ln both sides: 2x + 1 = ln1 -> 2x = ln1 -1 => x = (ln1 - 1)/2

OpenStudy (earthcitizen):

hint! ln1=0